### v1

@@ 5 On the semismooth property of solution maps to inclusions
We claim that there is some neighborhood $U\subset\Omega$ of $\bar{x}$ and some real $\gamma>0$ such that for all $x\in U$ and all $(z^{*},x^{*},y^{*})\in\mathrm{gph}\,D^{*}F(x,\sigma(x),0)$ one has
$\|y^{*}\|\geq\frac{1}{\gamma}\|(x^{*},z^{*})\|.$
Assume on the contrary that there are sequences $x_{k}\to\bar{x}$ and $(z_{k}^{*},x_{k}^{*},y_{k}^{*})\in\mathrm{gph}\,D^{*}F(x_{k},\sigma(x_{k}),0)$ with $\|(x^{*}_{k},z^{*}_{k})\|=1$ and $y_{k}^{*}\to 0$ as $k\to\infty$. By possibly passing to a subsequence we can assume that $(x_{k}^{*},z_{k}^{*})$ converges to some $(x^{*},z^{*})$ with $\|(x^{*},z^{*})\|=1$. It follows that $(z^{*},x^{*},0)\in\mathrm{gph}\,D^{*}F(\bar{x},\sigma(\bar{x}),0)$ contradicting (11). Hence our claim holds true. Consequently, there holds
@@
for every $x\in U$. By possibly shrinking $U$ we may assume that $U$ is an open, bounded and convex neighborhood of $\bar{x}$. By invoking [21, Corollary 4.60] we obtain that $\sigma$ is Lipschitz near every $x\in U$ with
@@
Thus, by [28, Theorem 9.2], we obtain that $\sigma$ is Lipschitz continuous on $U$ with constant $\gamma$. Hence, by virtue of (9) we have
$D^{*}\sigma(x)(y^{*})\subset\Sigma(x,\sigma(x),0)\ \forall x\in U\forall y^{*}\in\mathbb{R}^{m}.$ (12)
@@
Hence, for every $u\in\tilde{U}$ and every ${u^{*}}^{T}\in\Psi(u)$ we have
@@
verifying the local boundedness of $\Psi$ at $x$. Further, since $\mathrm{dom}\,\Phi$ is a neighborhood of $(x,\sigma(x))$ and taking into account (12), we obtain that $\mathrm{dom}\,\Psi$ is a neighborhood of $x$. Finally, let $\epsilon>0$ be given and consider $\delta>0$ such that for all $x^{\prime}\in U$ with $(x^{\prime},\sigma(x^{\prime}),0)\in{\cal B}_{\delta}(x,\sigma(x),0)$ and all $(z^{*},x^{*},y^{*})\in\mathrm{gph}\,D^{*}F(x^{\prime},\sigma(x^{\prime}),0)$ one has
$|\langle z^{*},0-0\rangle-\langle x^{*},x^{\prime}-x\rangle-\langle y^{*},\sigma(x^{\prime})-\sigma(x)\rangle|\leq\epsilon\|z^{*},x^{*},y^{*}\|\|(x^{\prime}-x,\sigma(x^{\prime})-\sigma(x),0-0)\|$
according to the semismoothness property of $F$ at $(x,\sigma(x),0)$. By possibly decreasing $\delta$ we may also assume that for all $x^{\prime}\in\tilde{U}$ with $(x^{\prime},\sigma(x^{\prime}))\in{\cal B}_{\delta}(x,\sigma(x))$ and all $(g_{x}^{T},g_{y}^{T})\in\Phi(x^{\prime},\sigma(x^{\prime}))$ there holds
@@
Choose $\delta^{\prime}\in(0,\delta/\sqrt{1+\gamma^{2}})$ such that ${\cal B}_{\delta^{\prime}}(x)\subset U$ and consider $x^{\prime}\in{\cal B}_{\delta^{\prime}}$ and ${u^{*}}^{T}\in\Psi(x^{\prime})$ together with $(g_{x}^{T},g_{y}^{T})\in\Phi(x^{\prime},\sigma(x^{\prime}))$ and $x^{*}\in\Sigma(x^{\prime},\sigma(x^{\prime}))(g_{y})$ such that $u^{*}=g_{x}^{T}+{x^{*}}^{T}$. Further choose $z^{*}$ such that $(z^{*},x^{*},-g_{y})\in\mathrm{gph}\,D^{*}F(x^{\prime},\sigma(x^{\prime}),0)$ according to the definition of $\Sigma$. Then $(x^{\prime},\sigma(x^{\prime})\in{\cal B}_{\delta}(x,\sigma(x))$ and consequently
@@
verifying that $\theta$ is semismooth with respect to $\Psi$ at $x$. This completes the proof. ∎
@@
In the setting of Theorem 5.3, consider a point $x\in U$ where $F$ is semismooth at $(x,\sigma(x),0)$ and consider the linear functions $\varphi_{i}:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$, $\varphi_{i}(x,y)=y_{i}$, $i=1,\ldots,m$, together with mappings $\tilde{\Psi}_{i}:U\rightrightarrows\mathbb{R}^{1\times n}$ given by $\tilde{\Psi}_{i}(u):=\{{u^{*}}^{T}\,\big{|}\,u^{*}\in\Sigma(u,\sigma(u))(-e_{i})$, where $e_{i}$ denotes the $i$-th unit vector in $\mathbb{R}^{m}$. It follows from Theorem 5.3 that $\sigma_{i}(u)=\varphi_{i}(u,\sigma(u))$ is semismooth at $x$ with respect to $\tilde{\Psi}_{i}$, $i=1,\ldots,m$. Hence $\sigma$ is also semismooth with respect to the mapping $\tilde{\Psi}$, which assigns to each $u\in U$ the set of matrices $A$ whose $i$-th row is an element from $\tilde{\Psi}_{i}(u)$.
@@
In the setting of Theorem 5.1, assume that we are given an open set $\Omega$ and a continuous mapping $\sigma:\Omega\to\mathbb{R}^{m}$ such that for every $x\in\Omega$ the mapping $F$ is semismooth at $(x,\sigma(x),0)$, the qualification condition
$(x^{*},0)\in D^{*}F(x,\sigma(x),0)(z^{*})\ \Rightarrow z^{*}=0,x^{*}=0$
holds and there exists a neighborhood $U\times V$ of $(x,\sigma(x))$ such that $\mathrm{gph}\,S\cap(U\times V)=\mathrm{gph}\,\sigma\cap(U\times V)$. Then the following statements hold:
  1. 1.
    $\sigma$ is locally Lipschitz on $\Omega$.
  2. 2.
    $\sigma$ is semismooth on $\Omega$ with respect to some mapping $\tilde{\Psi}:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n\times m}$, which can be constructed analogously to Remark 5.4.
  3. 3.
    Consider a function $\varphi:D\to\mathbb{R}$, where $D$ is an open subset of $\mathbb{R}^{n}\times\mathbb{R}^{m}$ containing $\mathrm{gph}\,\sigma$ and assume that $\varphi$ is semismooth on $\mathrm{gph}\,\sigma$ with respect to some mapping $\Phi:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{1\times(n+m)}$. Then the mapping $\theta:\Omega\to\mathbb{R}$ given by $\theta(x)=\varphi(x,\sigma(x))$ is semismooth on $\Omega$ with respect to the mapping $\Psi:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{1\times n}$ given by
    $\Psi(x):=\begin{cases}\{g_{x}^{T}+{x^{*}}^{T}\,\big{|}\,x^{*}\in\Sigma(x,\sigma(x))(g_{y}),(g_{x}^{T},g_{y}^{T})\in\Phi(x,\sigma(x))\}&\mbox{if $x\in\Omega$,}\\ \emptyset&\mbox{else.}\end{cases}$
Statement 3 enables us to apply the ImP approach to the numerical solution of a class of problems (2) with nonsmooth objectives $\varphi$ via a suitable bundle method. Apart from the case when $\sigma$ is a single-valued and Lipschitzian localization of $S$ for all $x\in\Omega$, however, the construction of $\tilde{\Psi}$ according to Remark 5.4 is rather complicated. We will now show that the theory considerably simplifies in the case of SCD mappings.
For an SCD mapping $F:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{m}$, the subspaces $L^{*}\in{\mathcal{S}}^{*}F(x,y,z)$ belong to ${\cal Z}_{m,n+m}$ and therefore $L^{*}\subset\mathbb{R}^{m+n+m}$. By identifying $\mathbb{R}^{m+n+m}$ with $\mathbb{R}^{m}\times\mathbb{R}^{n}\times\mathbb{R}^{m}$, we denote the elements of $L^{*}$ by $(z^{*},x^{*},y^{*})\in\mathbb{R}^{m}\times\mathbb{R}^{n}\times\mathbb{R}^{m}$. We start our analysis with the following result:
@@
For every subspace $L^{*}\in{\cal Z}_{m,{n+m}}$ satisfying
$(z^{*},x^{*},0)\in L^{*}\Rightarrow z^{*}=0,\ x^{*}=0$ (13)
@@
$L^{*}={\rm rge\;}(Z_{L^{*}},X_{L^{*}},I)=\{(Z_{L^{*}}p,X_{L^{*}}p,p)\,\big{|}\,p\in\mathbb{R}^{m}\}.$ (14)
In addition, for the norm
$\kappa(L^{*}):=\big{\|}\begin{pmatrix}Z_{L}^{*}\\ X_{L^{*}}\end{pmatrix}\big{\|}$
there holds
$\kappa(L^{*})=\sup\{\|(z^{*},x^{*})\|\,\big{|}\,(z^{*},x^{*},y^{*})\in L^{*},\|y^{*}\|\leq 1\}.$
@@
We claim that condition (13) ensures the nonsingularity of the matrix $B_{y^{*}}$. Assume on the contrary that there is some $0\not=p\in\mathbb{R}^{m}$ satisfying $B_{y^{*}}p=0$. Then we have $Bp=(B_{z^{*}}p,B_{x^{*}}p,0)\in L^{*}$ and consequently $B_{z^{*}}p=0$, $B_{x^{*}}p=0$ due to (13). Thus $Bp=0$, which is not possible since the colums of $B$ are a basis for $L^{*}$ and are therefore linearly independent. Hence the matrix $B_{y^{*}}$ is nonsingular and the columns of the matrix $BB_{y^{*}}^{-1}$ form another basis for $L^{*}$. Representation (14) now follows with
$Z_{L^{*}}=B_{z^{*}}B_{y^{*}}^{-1},\ X_{L^{*}}=B_{x^{*}}B_{y^{*}}^{-1}$
and it is easy to see that it is unique. Finally,
@@
Definition 5.7.
We denote by ${\cal Z}_{m,n+m}^{\rm reg}$ the collection of all subspaces $L^{*}\in{\cal Z}_{m,n+m}$ satisfying (13).
Remark 5.8.
From (14) it easily follows that for any $L^{*}\in{\cal Z}_{m,n+m}^{\mathrm{reg}\,}$ the corresponding primal subspace $L\in{\cal Z}_{n+m,m}$ has the representation
$L=\{(p_{x},X_{L^{*}}^{T}p_{x}+Z_{L^{*}}^{T}p_{z},p_{z})\,\big{|}\,(p_{x},p_{z})\in\mathbb{R}^{n}\times\mathbb{R}^{m}\}.$
In the following statements we collect some useful relationships concerning the subspaces $L^{*}$ from ${\cal Z}_{m,n+m}^{\mathrm{reg}\,}$ and the moduli $\kappa(L^{*})$ and ${\rm scd\,reg\,}F(x,y,z)$ (introduced below).
@@
    $(z^{*},x^{*},0)\in(\bigcup{\mathcal{S}}^{*}F)(x,y,z)\Rightarrow z^{*}=0,\ x^{*}=0.$ (15)
@@
The implication (i)$\Rightarrow$(ii) follows immediately from Proposition 5.6. We prove the implication (ii)$\Rightarrow$(iii). Assuming that (iii) does not hold there are sequences $(z_{k}^{*},x_{k}^{*},y_{k}^{*})\in L_{k}^{*}\subset(\bigcup{\mathcal{S}}^{*}F)(x,y,z)$ verifying $\|y_{k}^{*}\|\leq 1$ for all $k$ and $\|(z_{k}^{*},x_{k}^{*})\|\to\infty$ as $k\to\infty$. Since $L^{*}_{k}$ is a subspace, we have $(\tilde{z}^{*}_{k},\tilde{x}^{*}_{k},\tilde{y}^{*}_{k}):=(z_{k}^{*},x_{k}^{*},y_{k}^{*})/\|(z_{k}^{*},x_{k}^{*},y_{k}^{*})\|\in L_{k}^{*}$ for all $k$. Taking into account that the metric space ${\cal Z}_{m,n+m}$ is compact, by possibly passing to a subsequence we can assume that $L_{k}^{*}$ converges to some $L^{*}\in{\cal Z}_{m,n+m}$ and $(\tilde{z}^{*}_{k},\tilde{x}_{*}^{k},\tilde{y}^{*}_{k})$ converges to some $(\tilde{z}^{*},\tilde{x}^{*},\tilde{y}^{*})\in L$ with $\|(\tilde{z}^{*},\tilde{x}^{*},\tilde{y}^{*})\|=1$. Since ${\mathcal{S}}^{*}F(x,y,z)$ is closed in ${\cal Z}_{m,n+m}$, there holds $L^{*}\in{\mathcal{S}}^{*}F(x,y,z)$ and because of $\tilde{y}_{k}^{*}/\|z_{k}^{*},x_{k}^{*},y_{k}^{*}\|\to 0$ we have $\tilde{y}^{*}=0$. Hence $(\tilde{z}^{*},\tilde{x}^{*},0)\in L^{*}\in{\mathcal{S}}^{*}F(x,y,z)$ and $\|\tilde{z}^{*},\tilde{x}^{*})\|=1$ showing $L^{*}\not\in{\cal Z}_{m,n+m}^{\rm reg}$. This proves the implication (ii)$\Rightarrow$(iii). The implication (iii)$\Rightarrow$(i) holds trivially. Finally, formula (16) is a consequence of the definitions of ${\rm scd\,reg}F$, $(\bigcup{\mathcal{S}}^{*}F)$ and $\kappa(L^{*})$. ∎
@@
$\limsup_{(x^{\prime},y^{\prime},z^{\prime}){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(x,y,z)}{\rm scd\,reg\,}F(x^{\prime},y^{\prime},z^{\prime})\leq{\rm scd\,reg\,}F(x,y,z).$ (17)
Proof.
If $\gamma:={\rm scd\,reg\,}F(x,y,z)=\infty$ there is nothing to prove. Thus we can suppose that $\gamma<\infty$ and consequently, by Lemma 5.9, condition (15) is fulfilled. Assume now on the contrary that the inequality (17) does not hold and we can find sequences $(x_{k},y_{k},z_{k})\to(x,y,z)$ and $(z_{k}^{*},x_{k}^{*},y_{k}^{*})\in L_{k}^{*}\in{\mathcal{S}}^{*}F(x_{k},y_{k},z_{k})$ with $\|y_{k}\|^{*}\leq 1$ and $\limsup_{k\to\infty}\|(z^{*}_{k},x^{*}_{k})\|=:\bar{\gamma}>\gamma$. If $\bar{\gamma}$ is finite, by possibly passing to a subsequence we can assume that $(z_{k}^{*},x_{k}^{*},y_{k}^{*})$ converges to some $(\tilde{z}^{*},\tilde{x}^{*},\tilde{y}^{*})$ with $\bar{\gamma}=\|(\tilde{z}^{*},\tilde{x}^{*})\|$. By possibly passing to a subsequence once more we can assume that $L_{k}^{*}$ converges to some $L^{*}\in{\mathcal{S}}^{*}F(x,y,z)$. Since $(\tilde{z}^{*},\tilde{x}^{*},\tilde{y}^{*})\in L^{*}$ and $\|\tilde{y}^{*}\|\leq 1$, we obtain the contradiction $\gamma<\bar{\gamma}=\|(\tilde{z}^{*},\tilde{x}^{*})\|\leq\kappa(L^{*})\leq\gamma$. If $\bar{\gamma}=\infty$ we can employ the same arguments already used in the proof of Lemma 5.9 to obtain a contradiction. ∎
The next statement pertains to the $\Psi$-ss property of a single-valued continuous selection of (31) in the SCD case.
@@
Consider a mapping $F:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{m}$ and a point $(\bar{x},\bar{y})$ with $0\in F(\bar{x},\bar{y})$. Assume that $F$ has the SCD property around $(\bar{x},\bar{y},0)$ and ${\rm scd\,reg\,}F(\bar{x},\bar{y},0)<\infty$. Further suppose that there exists a neighborhood $\Omega$ of $\bar{x}$ and a continuous function $\sigma:\Omega\to\mathbb{R}^{m}$ verifying $\sigma(\bar{x})=\bar{y}$ and $0\in F(x,\sigma(x))$, $x\in\Omega$. Then there exists an open neighborhood $U\subset\Omega$ of $\bar{x}$ such that for every $x\in U$ there holds ${\rm scd\,reg\,}F(x,\sigma(x),0)<\infty$ and $\sigma$ is semismooth at $x$ with respect to the mapping $\Psi:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ given by
$\Psi(x):=\begin{cases}\{-X_{L^{*}}^{T}\,\big{|}\,L^{*}\in{\mathcal{S}}^{*}F(x,\sigma(x),0)\}&\mbox{if $x\in U$},\\ \emptyset&\mbox{else,}\end{cases}$

### v2

@@ 5 On the semismooth property of solution maps to inclusions
We claim that there is some neighborhood $U\subset\Omega$ of $\bar{x}$ and some real $\gamma>0$ such that for all $x\in U$ and all $(z^{*},x^{*},y^{*})\in\mathrm{gph}\,D^{*}F(x,\sigma(x),0)$ one has $\|y^{*}\|\geq\|(x^{*},z^{*})\|/\gamma.$ Assume on the contrary that there are sequences $x_{k}\to\bar{x}$ and $(z_{k}^{*},x_{k}^{*},y_{k}^{*})\in\mathrm{gph}\,D^{*}F(x_{k},\sigma(x_{k}),0)$ with $\|(x^{*}_{k},z^{*}_{k})\|=1$ and $y_{k}^{*}\to 0$ as $k\to\infty$. By possibly passing to a subsequence we can assume that $(x_{k}^{*},z_{k}^{*})$ converges to some $(x^{*},z^{*})$ with $\|(x^{*},z^{*})\|=1$. It follows that $(z^{*},x^{*},0)\in\mathrm{gph}\,D^{*}F(\bar{x},\sigma(\bar{x}),0)$ contradicting (5.16). Hence our claim holds true. Consequently, there holds
@@
for every $x\in U$. By possibly shrinking $U$ we may assume that $U$ is an open, bounded and convex neighborhood of $\bar{x}$. By invoking [28, Corollary 4.60] we obtain that $\sigma$ is Lipschitz near every $x\in U$ with
@@
Thus, by [35, Theorem 9.2], we obtain that $\sigma$ is Lipschitz continuous on $U$ with constant $\gamma$. Hence, by virtue of (5.14) we have
$D^{*}\sigma(x)(y^{*})\subset\Sigma(x,\sigma(x),0)\ \forall x\in U\ \forall y^{*}\in\mathbb{R}^{m}.$ (5.17)
@@
Hence, for every $u\in\tilde{U}$ and every ${u^{*}}^{T}\in\Theta(u)$ we have
@@
verifying the local boundedness of $\Theta$ at $x$. Further, since $\mathrm{dom}\,\Theta$ is a neighborhood of $(x,\sigma(x))$ and taking into account (5.17), we obtain that $\mathrm{dom}\,\Theta$ is a neighborhood of $x$. Finally, let $\epsilon>0$ be given and consider $\delta>0$ such that for all $x^{\prime}\in U$ with $(x^{\prime},\sigma(x^{\prime}),0)\in{\cal B}_{\delta}(x,\sigma(x),0)$ and all $(z^{*},x^{*},y^{*})\in\mathrm{gph}\,D^{*}F(x^{\prime},\sigma(x^{\prime}),0)$ one has
$|\langle z^{*},0-0\rangle-\langle x^{*},x^{\prime}-x\rangle-\langle y^{*},\sigma(x^{\prime})-\sigma(x)\rangle|\leq\epsilon\|(z^{*},x^{*},y^{*})\|\|(x^{\prime}-x,\sigma(x^{\prime})-\sigma(x),0-0)\|$
due to the semismoothness property of $F$ at $(x,\sigma(x),0)$. By possibly decreasing $\delta$ we may also assume that for all $x^{\prime}\in\tilde{U}$ with $(x^{\prime},\sigma(x^{\prime}))\in{\cal B}_{\delta}(x,\sigma(x))$ and all $(g_{x}^{T},g_{y}^{T})\in\Phi(x^{\prime},\sigma(x^{\prime}))$ there holds
@@
Choose $\delta^{\prime}\in(0,\delta/\sqrt{1+\gamma^{2}})$ such that ${\cal B}_{\delta^{\prime}}(x)\subset U$ and consider $x^{\prime}\in{\cal B}_{\delta^{\prime}}(x)$ and ${u^{*}}^{T}\in\Theta(x^{\prime})$ together with $(g_{x}^{T},g_{y}^{T})\in\Phi(x^{\prime},\sigma(x^{\prime}))$ and $x^{*}\in\Sigma(x^{\prime},\sigma(x^{\prime}))(g_{y})$ such that $u^{*}=g_{x}^{T}+{x^{*}}^{T}$. Further choose $z^{*}$ such that $(z^{*},x^{*},-g_{y})\in\mathrm{gph}\,D^{*}F(x^{\prime},\sigma(x^{\prime}),0)$ according to the definition of $\Sigma$. Then $(x^{\prime},\sigma(x^{\prime}))\in{\cal B}_{\delta}(x,\sigma(x))$ and consequently
@@
verifying that $\theta$ is semismooth with respect to $\Theta$ at $x$. This completes the proof. ∎
@@
In the setting of Theorem 5.3, consider a point $x\in U$ where $F$ is semismooth at $(x,\sigma(x),0)$ and consider the linear functions $\beta_{i}:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$, $\beta_{i}(x,y)=y_{i}$, $i=1,\ldots,m$, together with mappings $\Psi_{i}:U\rightrightarrows\mathbb{R}^{1\times n}$ given by $\Psi_{i}(u):=\{{u^{*}}^{T}\,\big{|}\,u^{*}\in\Sigma(u,\sigma(u))(-e_{i})\}$, where $e_{i}$ denotes the $i$-th unit vector in $\mathbb{R}^{m}$. It follows from Theorem 5.3 that $\sigma_{i}(u)=\beta_{i}(u,\sigma(u))$ is semismooth at $x$ with respect to $\Psi_{i}$, $i=1,\ldots,m$. Hence, $\sigma$ is also semismooth with respect to the mapping $\Psi$, which assigns to each $u\in U$ the set of matrices whose $i$-th row is an element from $\Psi_{i}(u)$ as in (3.10).
@@
In the setting of Theorem 5.1, assume that we are given an open set $\Omega$ and a continuous mapping $\sigma:\Omega\to\mathbb{R}^{m}$ such that for every $x\in\Omega$ the mapping $F$ is semismooth at $(x,\sigma(x),0)$, the qualification condition (5.13) holds and there exists a neighborhood $U\times V$ of $(x,\sigma(x))$ such that $\mathrm{gph}\,S\cap(U\times V)=\mathrm{gph}\,\sigma\cap(U\times V)$.
Then $\sigma$ is semismooth on $\Omega$ with respect to some mapping $\Psi:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n\times m}$, which can be constructed analogously to Remark 5.4. Further, for every function $\varphi:D\to\mathbb{R}$, where $D$ is an open subset of $\mathbb{R}^{n}\times\mathbb{R}^{m}$ containing $\mathrm{gph}\,\sigma$, which is semismooth on $\mathrm{gph}\,\sigma$ with respect to some mapping $\Phi:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{1\times(n+m)}$, there holds that the mapping $\theta:\Omega\to\mathbb{R}$ given by $\theta(x)=\varphi(x,\sigma(x))$ is semismooth on $\Omega$ with respect to the mapping $\Theta:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{1\times n}$ given by
$\Theta(x):=\{g_{x}^{T}+{x^{*}}^{T}\,\big{|}\,x^{*}\in\Sigma(x,\sigma(x))(g_{y}),(g_{x}^{T},g_{y}^{T})\in\Phi(x,\sigma(x))\},\ x\in\mathrm{dom}\,\Theta:=\Omega.$
This corollary enables us to apply the ImP approach to the numerical solution of a class of problems (1.3) with nonsmooth objectives $\varphi$ via a suitable bundle method. Apart from the case when $\sigma$ is a single-valued and Lipschitz localization of $S$ for all $x\in\Omega$, however, the construction of $\Psi$ according to Remark 5.4 is rather complicated. We will now show that the theory considerably simplifies in the case of SCD mappings.
For an SCD mapping $F:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{m}$, the subspaces $L^{*}\in{\mathcal{S}}^{*}F(x,y,z)$ belong to ${\cal Z}_{m,n+m}$ and therefore $L^{*}\subset\mathbb{R}^{m+n+m}$. By identifying $\mathbb{R}^{m+n+m}$ with $\mathbb{R}^{m}\times\mathbb{R}^{n}\times\mathbb{R}^{m}$, we denote the elements of $L^{*}$ by $(z^{*},x^{*},y^{*})\in\mathbb{R}^{m}\times\mathbb{R}^{n}\times\mathbb{R}^{m}$. We start our analysis with the following auxiliary result:
@@
For every subspace $L^{*}\in{\cal Z}_{m,{n+m}}$ satisfying the implication
$(z^{*},x^{*},0)\in L^{*}\Rightarrow z^{*}=0,\ x^{*}=0$ (5.18)
@@
$L^{*}={\rm rge\;}(Z_{L^{*}},X_{L^{*}},I)=\{(Z_{L^{*}}p,X_{L^{*}}p,p)\,\big{|}\,p\in\mathbb{R}^{m}\}.$ (5.19)
In addition, there holds
$\kappa(L^{*}):=\big{\|}\begin{pmatrix}Z_{L}^{*}\\ X_{L^{*}}\end{pmatrix}\big{\|}=\sup\{\|(z^{*},x^{*})\|\,\big{|}\,(z^{*},x^{*},y^{*})\in L^{*},\|y^{*}\|\leq 1\}.$
@@
We claim that condition (5.18) ensures the nonsingularity of the matrix $B_{y^{*}}$. Assume on the contrary that there is some $0\not=p\in\mathbb{R}^{m}$ satisfying $B_{y^{*}}p=0$. Then we have $Bp=(B_{z^{*}}p,B_{x^{*}}p,0)\in L^{*}$ and consequently $B_{z^{*}}p=0$, $B_{x^{*}}p=0$ due to (5.18). Thus $Bp=0$, which is not possible since the columns of $B$ form a basis for $L^{*}$ and are therefore linearly independent. Hence the matrix $B_{y^{*}}$ is nonsingular and the columns of the matrix $BB_{y^{*}}^{-1}$ form another basis for $L^{*}$. Representation (5.19) now follows with $Z_{L^{*}}=B_{z^{*}}B_{y^{*}}^{-1},\ X_{L^{*}}=B_{x^{*}}B_{y^{*}}^{-1}$ and it is easy to see that it is unique. Finally,
@@
Let us denote by ${\cal Z}_{m,n+m}^{\rm reg}$ the collection of all subspaces $L^{*}\in{\cal Z}_{m,n+m}$ satisfying (5.18). In the following statements we collect some useful relations concerning the subspaces $L^{*}$ from ${\cal Z}_{m,n+m}^{\mathrm{reg}\,}$ and the moduli $\kappa(L^{*})$ and ${\rm scd\,reg\,}F(x,y,z)$ (introduced below).
@@
    $(z^{*},x^{*},0)\in(\bigcup{\mathcal{S}}^{*}F)(x,y,z)\Rightarrow z^{*}=0,\ x^{*}=0.$
@@
The implication (i)$\Rightarrow$(ii) follows immediately from Proposition 5.6. We prove the implication (ii)$\Rightarrow$(iii). Assuming that (iii) does not hold there are sequences $(z_{k}^{*},x_{k}^{*},y_{k}^{*})\in L_{k}^{*}\subset(\bigcup{\mathcal{S}}^{*}F)(x,y,z)$ with $\|y_{k}^{*}\|\leq 1$ for all $k$ and $\|(z_{k}^{*},x_{k}^{*})\|\to\infty$ as $k\to\infty$. Since $L^{*}_{k}$ is a subspace, we have $(\tilde{z}^{*}_{k},\tilde{x}^{*}_{k},\tilde{y}^{*}_{k}):=(z_{k}^{*},x_{k}^{*},y_{k}^{*})/\|(z_{k}^{*},x_{k}^{*},y_{k}^{*})\|\in L_{k}^{*}$ for all $k$. Taking into account that the metric space ${\cal Z}_{m,n+m}$ is compact, by possibly passing to a subsequence we can assume that $L_{k}^{*}$ converges to some $L^{*}\in{\cal Z}_{m,n+m}$ and $(\tilde{z}^{*}_{k},\tilde{x}^{*}_{k},\tilde{y}^{*}_{k})$ converges to some $(\tilde{z}^{*},\tilde{x}^{*},\tilde{y}^{*})\in L^{*}$ with $\|(\tilde{z}^{*},\tilde{x}^{*},\tilde{y}^{*})\|=1$. Since ${\mathcal{S}}^{*}F(x,y,z)$ is closed in ${\cal Z}_{m,n+m}$, there holds $L^{*}\in{\mathcal{S}}^{*}F(x,y,z)$ and because of $\tilde{y}_{k}^{*}/\|(z_{k}^{*},x_{k}^{*},y_{k}^{*})\|\to 0$ we have $\tilde{y}^{*}=0$. Hence $(\tilde{z}^{*},\tilde{x}^{*},0)\in L^{*}\in{\mathcal{S}}^{*}F(x,y,z)$ and $\|(\tilde{z}^{*},\tilde{x}^{*})\|=1$ showing $L^{*}\not\in{\cal Z}_{m,n+m}^{\rm reg}$. This proves the implication (ii)$\Rightarrow$(iii). The implication (iii)$\Rightarrow$(i) holds trivially. Finally, formula (5.20) is a consequence of the definitions of ${\rm scd\,reg}F$, $(\bigcup{\mathcal{S}}^{*}F)$ and $\kappa(L^{*})$. ∎
By following the proof of [12, Proposition 4.8] one can easily show the following result.
@@
$\limsup_{(x^{\prime},y^{\prime},z^{\prime}){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(x,y,z)}{\rm scd\,reg\,}F(x^{\prime},y^{\prime},z^{\prime})\leq{\rm scd\,reg\,}F(x,y,z).$
The next statement pertains to the $\Psi$-ss property of a single-valued continuous selection of (5.12) in the SCD case.
@@
Consider a mapping $F:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{m}$ and a point $(\bar{x},\bar{y})$ with $0\in F(\bar{x},\bar{y})$. Assume that $F$ has the SCD property around $(\bar{x},\bar{y},0)$ and ${\rm scd\,reg\,}F(\bar{x},\bar{y},0)<\infty$. Further suppose that there exists a neighborhood $\Omega$ of $\bar{x}$ and a continuous function $\sigma:\Omega\to\mathbb{R}^{m}$ with $\sigma(\bar{x})=\bar{y}$ and $0\in F(x,\sigma(x))$, $x\in\Omega$. Then there exists an open neighborhood $U\subset\Omega$ of $\bar{x}$ such that for every $x\in U$ there holds ${\rm scd\,reg\,}F(x,\sigma(x),0)<\infty$ and $\sigma$ is semismooth at $x$ with respect to the mapping $\Psi:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m\times n}$ given by
$\Psi(x):=\{-X_{L^{*}}^{T}\,\big{|}\,L^{*}\in{\mathcal{S}}^{*}F(x,\sigma(x),0)\},\ x\in\mathrm{dom}\,\Psi:=U,$